# ISAC channel simulator GBSM

### v1

@@ 2.2.1 RCS polarization
The scattering characteristics of the target are related to the polarization of the incident field. When the antenna is linearly polarized, the linearly polarized wave incident on the target surface can be decomposed into vertical and horizontal polarization components. The polarization components of the target’s scattered field received by the receiving antenna can be expressed as:
@@ 2.2.2 Swerling modeling
For complex targets, when the spacing between scattering elements is comparable to the wavelength, changes in the observation direction or operating wavelength alter the phase of each scattering channel received at the input of the receiver, thereby modifying the resultant vector and producing fluctuating target signals. The size of the Radar Cross Section (RCS) directly affects radar detection performance, and in engineering calculations, RCS is often treated as a constant, as shown by the average values in Table 1. For targets in motion, the cross-sectional area fluctuates with the viewing angle. To accurately characterize the fluctuations of the radar cross section, its probability density function $p(\sigma)$ must be provided. The Swerling fluctuation model is a classic and widely used model in radar applications. It categorizes fluctuations into four types based on the correlation time and fluctuation rate:
@@
where $\overline{\sigma}$ is the average value of RCS, $\sigma$ follows an exponential distribution and $\sigma>0$.The probability distribution functions of Swerling 1 and Swerling 2 are Identical, but Swerling 1 Exhibits Slow Fluctuations While Swerling 2 Exhibits Fast Fluctuations.
@@
The probability distribution functions of Swerling 3 and Swerling 4 are Identical, in which Swerling 3 Exhibits Slow Fluctuations and Swerling 4 Exhibits Fast Fluctuations.The first and second types of fluctuation models are applicable to complex targets composed of a large number of approximately equal scatterers, such as an aircraft.The third and fourth types of fluctuation models are suitable for targets that consist of one large reflecting surface combined with many small scattering surfaces or for a single large scattering surface with small changes in aspect.
@@ 2.2.3 Measurement based modeling
Based on the extensive frequency, distance, and target type-dependent measurement analysis conducted by the team in [15], the following conclusions can be drawn regarding RCS modeling:
  • The RCS of the UAV is angle-independent and exhibits similar properties to omnidirectional scattering. In addition, the measurement results from mono-static mode and bi-static mode are found to be closely approximated.
  • The RCS model of the UAV is modeled as a normal distribution in both mono-static and bi-static sensing modes. The mean is used to determine the deterministic component and the variance is used to determine the randomly generated component.
  • The RCS of the human is angle-independent and exhibits similar properties to omnidirectional scattering. In addition, the measurement results from mono-static mode and bi-static mode are found to be closely approximated.
  • The RCS model of the human is modeled as a normal distribution in both mono-static and bi-static sensing modes. The mean is used to determine the deterministic component and the variance is used to determine the randomly generated component.
Different from human body and UAV, vehicle RCS have spiking characteristics and are not suitable for fitting with omnidirectional patterns. The RCS of humans and UAVs can be fitted using a log-normal distribution, whereby the mean $\mu$ and variance $\sigma$ are derived from empirical measurements for random generation. The RCS of vehicles exhibits four peaks in its angular distribution, necessitating further discussion on its modeling and simulation methods.
@@
This equation holds under the condition that the target is sufficiently smooth, without the presence of shadowing and backscatter. The bistatic RCS for a specific target requires support from measurement results and can be modeled similarly to an antenna radiation pattern.
@@ 3.1 Accuracy metric
The basic detection equation in the field of radar is as follows in 2. Here,$c$ is the speed of light with a value of $3\times 10^{8}m/s$ and $t_{R}$ represents the time difference between the transmitted pulse and the received echo pulse from the sensed target. The ranging accuracy can be qualitatively analyzed using the total differential expression of equation LABEL:eq_RDtheory. The range resolution $\Delta R$ is defined as the minimum distinguishable distance, which is related to the pulse width $t_{\tau}$ of the sensing signal. The maximum unambiguous range is determined by the pulse repetition interval $T_{R}$ and is expressed as follows:
@@
where $\phi$ represents the phase difference between different antenna elements that are separated by a distance $d$ at the receiver and $\theta$ is the angle of sensing target.The angle measurement accuracy can be expressed as
@@
from equation 22,$\theta_{max}=\arcsin{(\frac{\lambda}{2\pi})}$. The angle resolution $\Delta\theta$ is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array $\frac{2\pi}{N}$. according to LABEL:eq_ADresolution,$\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$ where $N$ is the number of antenna array. the basic principle of velocity measurement is based on the Doppler frequency. from the above equation,we obtain $v=\frac{\lambda f_{d}}{2}$.analogous to range resolution and angular resolution. Here, the Doppler resolution $\Delta f_{d}=\frac{1}{MT_{R}}$, in which $M$ is the number of transmitted pulses, and $T_{R}$ is the pulse repetition interval.
@@ 3.2 Detection metric
False alarm probability occurs when the noise at the receiver exceeds the decision threshold, leading to a false detection. This probability is related to the root mean square (RMS) value of the noise and the decision threshold, and is defined as follows.
@@ 4 Simulation Results
In the simulation experiments, the focus is primarily on the reliability of the ISAC channel cascading method and the RCS simulation results. To validate the channel cascading method, specific scenario parameters used in the dual-station sensing mode of a UAV scenario are presented in Table X. The statistical characteristics of the generated channels, including delay spread, horizontal angle of arrival (AoA) spread, horizontal angle of departure (AoD) spread, vertical AoA spread, and vertical AoD spread, are validated for the 5G standard channel model, full convolution method, and the proposed parameter convolution method. The number of simulation runs is set to 1000 to collect statistics on the generated channel’s spread values. For the validation of the RCS simulation results, the mean and variance of the RCS concerning the log-normal distribution are obtained by fitting the measured RCS results reported by the team in document [15]. The RCS is simulated based on the fitted distribution, and the simulated RCS results are compared against the measured data to evaluate the error.
Figure 3: Delay sperad of ISAC concatenated Channel
As shown in Fig. 3, the delay spread of the cascaded channels is always greater than that of the standard channel, which is attributed to the addition of the two delay components during the cascading process, resulting in a broader overall delay distribution function and an increase in the spread value statistics. The spread value under the parameter convolution method is smaller than that of the full convolution method, as the parameter convolution approach reduces complexity through cluster pruning, thereby eliminating low-power, high-delay clusters, which in turn reduces the spread value.
@@
Figure 5: RCS of UAV and Human in 5GHz 14GHz 28GHz respectively
Fig. 4 illustrates that the statistical values of the angle spread for the three channels closely align, especially in the horizontal angle spread dimension. This is due to the cascading scheme generating statistical parameters for each segment, while the mean and variance distributions remain consistent across simulations of the same scenario, leading to a convergence of spread values through multiple statistical averages. A deviation is observed in the ESD, where the spread value of the cascaded scheme exceeds that of the standard channel. This is because the cluster pruning retains high-power clusters, which are closely aligned with the angle of arrival. Consequently, after pruning and cascading, high-power clusters are duplicated, leading to an increase in the spread value. The spread value of the parameter convolution method is lower than that of the full convolution due to the reduction of clusters with larger angles after pruning. Fig. 5 presents the RCS simulation results for two target types, Human and UAV, at frequencies of 5 GHz, 14 GHz, and 28 GHz. The simulated results align well with the measured results reported in proposal x, validating the feasibility of the log-normal fitting scheme. Additionally, it confirms the differences in RCS distribution between integrated sensing and communication channels and traditional radar systems. This discrepancy arises because traditional radar RCS perception occurs at longer distances, whereas integrated sensing and communication services focus on closer ranges, leading to variations between the RCS distributions in radar applications and the measured RCS results in ISAC.
@@ 5 Conclusion
In this paper, we extended the 3D-GBSM 5G standard channel model, integrating the latest advancements in 3GPP ISAC standardization to enhance the sensing and communication integrated channel simulation capabilities for 6G. The new model supports both Bi-Static and Mono-Static sensing modes, accommodating four types of sensing targets: UAVs, vehicles, humans, and AGVs. It aligns with various options outlined in the 3GPP standardization progress, including modeling methods for RCS size and scale, polarization techniques, and the power normalization methods for target and background channels.
By employing a parameter cascading method, we preserved the statistical characteristics of the ISAC channel distribution while achieving a 93.75% reduction in average convolution complexity. The simulation database for target RCS was developed through the fitting of actual measurements related to corresponding target RCS. A comparative analysis demonstrated good consistency between the simulated RCS and the measured RCS.
This advancement accelerates the standardization of the 6G ISAC channel model and the evaluation of new algorithms, as we can now utilize existing 5G standard parameter tables along with the proposed cascading model and measured RCS data to simulate ISAC channels. The modified simulation framework is applicable to all future developments in ISAC channel model standardization, providing insights for the standardization simulation process and expediting simulation evaluation progress.

### v2

@@ 2.2.1 RCS polarization
The scattering characteristics of a target are influenced by the polarization of the incident field. When the antenna is linearly polarized, the incident linearly polarized wave on the target surface can be decomposed into vertical and horizontal polarization components. The polarization components of the scattered field from the target, as received by the receiving antenna, can be expressed as:
@@ 2.2.2 Swerling modeling
For complex targets, when the spacing between scattering elements is comparable to the wavelength, changes in the observation direction or operating wavelength alter the phase of each scattering channel received at the input of the receiver, thereby modifying the resultant vector and producing fluctuating target signals. The size of the Radar Cross Section (RCS) directly influences radar detection performance. In engineering calculations, RCS is frequently treated as a constant, typically represented by its average values. For moving targets, the RCS varies with the viewing angle. To accurately characterize these fluctuations, it is essential to provide its probability density function $p(\sigma)$. The Swerling fluctuation model is a classic and widely utilized framework in radar applications. It categorizes fluctuations into four distinct types based on correlation time and fluctuation rate:
@@
where $\overline{\sigma}$ is the average value of RCS, $\sigma$ follows an exponential distribution and $\sigma>0$.The probability distribution functions of Swerling 1 and Swerling 2 are identical; however, Swerling 1 exhibits slow fluctuations, while Swerling 2 demonstrates fast fluctuations.
@@
The probability distribution functions of Swerling 3 and Swerling 4 are identical; however, Swerling 3 exhibits slow fluctuations, while Swerling 4 demonstrates fast fluctuations. The first and second types of fluctuation models apply to complex targets composed of a large number of approximately equal scatterers, such as aircraft. The third and fourth types of fluctuation models are appropriate for targets characterized by a large reflecting surface in conjunction with numerous small scattering surfaces, or for a singular large scattering surface exhibiting minor variations in aspect.
@@ 2.2.3 Measurement based modeling
Drawing upon the comprehensive analysis of measurements that depend on frequency, distance, and target type conducted by the team in [15], the following conclusions regarding RCS modeling can be established:
  • The RCS of the UAV is independent of angle and demonstrates characteristics akin to omnidirectional scattering. Furthermore, the measurement results obtained from both mono-static and bi-static modes closely approximate each other.
  • The RCS model for the UAV is represented as a normal distribution in both mono-static and bi-static sensing modes. The mean defines the deterministic component, while the variance characterizes the stochastic component.
  • The RCS of a human is angle-independent and displays characteristics akin to omnidirectional scattering. Furthermore, the measurement results obtained from mono-static and bi-static modes closely align with each other.
  • The RCS model for humans is represented as a normal distribution in both mono-static and bi-static sensing modes. The mean value characterizes the deterministic component, while the variance accounts for the stochastic component.
Different from the human body and UAV, vehicle RCS exhibit spiking characteristics that render them unsuitable for modeling with omnidirectional patterns. The RCS of humans and UAVs can be modeled using a log-normal distribution, in which the mean $\mu$ and variance $\sigma$ are determined from empirical measurements for random generation. displays four distinct peaks in its angular distribution, indicating the need for further exploration of its modeling and simulation methodologies.
@@
This equation is valid under the condition that the target is sufficiently smooth and free from shadowing and backscatter effects. necessitates validation through measurement results and can be modeled analogous to an antenna radiation pattern.
@@ 3.1 Accuracy metric
The fundamental detection equation in radar is presented in Table 2. Here,$c$ is the speed of light with a value of $3\times 10^{8}m/s$ and $t_{R}$ represents the time difference between the transmitted pulse and the received echo pulse from the sensed target. The ranging accuracy can be qualitatively analyzed using the total differential expression of equation 19. The range resolution $\Delta R$ is defined as the minimum distinguishable distance, which is related to the pulse width $t_{\tau}$ of the sensing signal. The maximum unambiguous range is determined by the pulse repetition interval $T_{R}$ and is expressed as follows:
@@
where $\phi$ represents the phase difference between different antenna elements that are separated by a distance $d$ at the receiver and $\theta$ is the angle of the sensing target. The angle measurement accuracy can be expressed as
@@
from equation 22,$\theta_{max}=\arcsin{(\frac{\lambda}{2\pi})}$. The angle resolution $\Delta\theta$ is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array $\frac{2\pi}{N}$. according to LABEL:eq_ADresolution,$\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$ where $N$ is the number of antenna arrays. the basic principle of velocity measurement is based on the Doppler frequency. from the above equation,we obtain $v=\frac{\lambda f_{d}}{2}$.analogous to range resolution and angular resolution. Here, the Doppler resolution $\Delta f_{d}=\frac{1}{MT_{R}}$, in which $M$ is the number of transmitted pulses, and $T_{R}$ is the pulse repetition interval.
@@ 3.2 Detection metric
The probability of false alarm arises when the noise at the receiver surpasses the decision threshold, resulting in erroneous detection. This probability is associated with the root mean square (RMS) value of the noise and the decision threshold, and it is defined as follows
@@ 4 Simulation Results
In the simulation experiments, the primary focus is on the reliability of the ISAC channel cascading method and the accuracy of the RCS simulation results. To validate the channel cascading method, specific scenario parameters utilized in the dual-station sensing mode for a UAV scenario. The statistical characteristics of the generated channels—specifically, delay spread, horizontal angle of arrival (AoA) spread, horizontal angle of departure (AoD) spread, vertical AoA spread, and vertical AoD spread—are validated against the 5G standard channel model, the full convolution method, and the proposed parameter convolution method. The number of simulation runs is established at 1000 to collect statistical data on the spread values of the generated channel. To validate the RCS simulation results, the mean and variance of the RCS, based on the log-normal distribution, are derived by fitting the measured RCS results reported in document [15]. The RCS is simulated using the fitted distribution, and the simulated RCS results are compared with the measured data to assess the error.
Figure 3: Delay spread of ISAC concatenated Channel
As illustrated in Fig. 3, the delay spread of the cascaded channels consistently exceeds that of the standard channel. This phenomenon is attributed to the accumulation of the two delay components during the cascading process, which results in a broader overall delay distribution function and an increase in the spread value statistics. The spread value obtained through the parameter convolution method is smaller than that achieved with the full convolution method. This reduction is due to the parameter convolution approach, which decreases complexity by pruning clusters, thus eliminating low-power, high-delay clusters and subsequently reducing the spread value.
@@
Figure 5: RCS of UAV and Human in 5GHz, 14GHz, and 28GHz respectively
Fig. 4 demonstrates that the statistical values of the angle spread for the three channels closely align, particularly in the dimension of horizontal angle spread. This phenomenon arises from the cascading scheme, which generates statistical parameters for each segment. The mean and variance distributions remain consistent across simulations of the same scenario, resulting in the convergence of spread values through multiple statistical averages. A deviation is observed in 4, where the spread value of the cascaded scheme exceeds that of the standard channel. This phenomenon arises from the cluster pruning process retaining high-power clusters that are closely aligned with the angle of arrival. Consequently, the duplication of high-power clusters following pruning and cascading results in an increase in the spread value. The spread value associated with the parameter convolution method is lower than that of the full convolution method, attributed to the reduction of clusters with larger angles following pruning. Figure 5 illustrates the RCS simulation results for two target types: human and UAV, at frequencies of 5 GHz, 14 GHz, and 28 GHz. The simulated results demonstrate a strong alignment with the measured results reported in reference [14], thereby validating the feasibility of the log-normal fitting scheme. Additionally, this finding confirms the differences in RCS distribution between ISAC channels and traditional radar systems. This discrepancy arises from the fact that traditional radar RCS perception typically occurs at greater distances, while ISAC services are concentrated on closer ranges. Consequently, this focus results in variations between the RCS distributions observed in radar applications and those measured in ISAC.
@@ 5 Conclusion
In this paper, we extend the 3D-GBSM 5G standard channel model by incorporating the latest advancements in 3GPP ISAC standardization, thereby enhancing the integrated sensing and communication channel simulation capabilities for 6G. The new model accommodates both Bi-Static and Mono-Static sensing modes, supporting four types of sensing targets: UAVs, vehicles, humans, and AGVs. This model aligns with various options outlined in the 3GPP standardization progress, including methods for modeling RCS size and scale, polarization techniques, and power normalization for target and background channels.
By employing a parameter cascading method, we preserved the statistical characteristics of the ISAC channel distribution while achieving a 93.75% reduction in average convolution complexity. The simulation database for target RCS was developed by fitting actual measurements corresponding to the target RCS. A comparative analysis demonstrated a high degree of consistency between the simulated RCS and the measured RCS.
This advancement accelerates the standardization of the 6G ISAC channel model and the evaluation of novel algorithms, as it enables the utilization of existing 5G standard parameter tables alongside the proposed cascading model and measured RCS data to simulate ISAC channels. The modified simulation framework is applicable to all future developments in ISAC channel model standardization, offering insights into the standardization simulation process and facilitating the progress of simulation evaluations.